# Crossed products by compact group actions with the weak tracial Rokhlin property

### v1

@@ 3. The weak tracial Rokhlin property with comparison
Without loss of generality, we may assume $F$ is $\alpha$-invariant. For the compact set $F$ and $\frac{\varepsilon}{2}$, by Lemma 3.14, we obtain a positive number $\delta>0$, a finite subset $K\subseteq G$, and a partition of unity $(f_{k})_{k\in K}\subseteq C(G)$ satisfying conditions (1)-(5) of Lemma 3.14. Set $S=\{f_{k},f_{k}^{\frac{1}{2}}\colon k\in K\}$. Choose an increasing sequence of finite dense subsets $S_{n}\subseteq C(G)$ such that $S\subseteq S_{n}$ for each $n\in\mathbb{N}$. Similarly, choose an increasing sequence of compact dense subsets $F_{n}\subseteq A$ such that $F\subseteq F_{n}$ for each $n\in\mathbb{N}$. For the finite set $S_{m}\subseteq C(G)$, the compact set $F_{m}\subseteq A$, $\frac{1}{m}$, $x$, and $y$, by Lemma 3.6, we obtain $d_{m}\in A^{\alpha}$ and a completely positive contractive equivariant $(S_{m},F_{m},\frac{1}{m})$-approximately central approximate order zero map $\psi_{m}\colon C(G)\to\overline{d_{m}Ad_{m}}$. As in Lemma 3.13, define a completely positive contractive order zero map $\psi\colon C(G)\to A_{\infty,\alpha}\cap A^{\prime}$. Identify $C(G)\otimes A$ with $C(G,A)$, and define a completely positive contractive map $\varphi\colon C(G,A)\to A_{\infty,\alpha}$ by $\varphi(f\otimes a)=\psi(f)a$ for all $f\in C(G)$ and $a\in A$ (note that the range of $\psi$ commutes with $A$ in $A_{\infty,\alpha}$). Clearly $\varphi$ is equivariant, where we equip $C(G,A)$ with the diagonal action $\gamma$ of $G$. For $m\in\mathbb{N}$, choose completely positive contractive maps $\varphi_{m}\colon C(G,A)\to A$ such that for all $\xi\in C(G,A)$,
$\pi_{A}((\varphi_{m}(\xi))_{m\in\mathbb{N}})=\varphi(\xi).$
Then for all $f\in C(G)$ and $a\in A$, $\limsup_{m\to\infty}\|\varphi_{m}(f\otimes a)-\psi_{m}(f)a\|=0$, and for all $g\in G$ and $\xi\in C(G,A)$, $\limsup_{m\to\infty}\|\alpha_{g}(\varphi_{m}(\xi))-\varphi_{m}(\gamma_{g}(\xi))\|=0$. We first verify condition (5).
Given $a\in F$ and $g\in G$, applying condition (5) of Lemma 3.14 in the last step and noting $\gamma_{g}(\varphi(a))=\varphi(a)$, we have
@@
$\leq$ ${}\limsup_{m\to\infty}\|\gamma_{g}(\sum_{k\in K}f_{k}\otimes\alpha_{k}(a))-\sum_{k\in K}f_{k}\otimes\alpha_{k}(a)\|\leq\frac{\varepsilon}{2}.$
Choose $m$ large enough such that $m>\frac{|K|^{2}}{\varepsilon}$ and
@@
For such $m$, condition (1) is obvious. Condition (2) is obvious due to the continuity of $\alpha$. Conditions (3) and (4) come from Lemma 3.14. Conditions (6)-(11) come from Lemma 3.6. ∎
We now present the most important theorem of this section, which is a key tool for transferring properties from the original algebra to the fixed point algebra. The proof references [18, Theorem 3.1].
@@
Without loss of generality, we may assume $\|a\|\leq 1$ for all $a\in F_{1}\cup F_{2}$. For the compact set $F=F_{1}\cup F_{2}$ (without loss of generality, assumed $\alpha$-invariant), $\varepsilon_{0}=\frac{\varepsilon}{7}$, $x\in A_{+}$ with $\|x\|=1$, and $y\in(A^{\alpha})_{+}\setminus\{0\}$, use Proposition 3.15 to obtain a positive number $\delta>0$, a finite subset $K\subseteq G$, a partition of unity $(f_{k})_{k\in K}\subseteq C(G)$, a contractive positive element $d\in A^{\alpha}$, and a completely positive contractive map $\psi\colon C(G)\to\overline{dAd}$ such that
@@
    If $g,g^{\prime}\in G$ satisfy $d(g,g^{\prime})<\delta$, then $\|\alpha_{g}(a)-\alpha_{g^{\prime}}(a)\|<\frac{\varepsilon_{0}}{|K|^{2}}$ for all $a\in F$.
@@
    $\|\alpha_{g}(\sum_{k\in K}\psi(f_{k})^{\frac{1}{2}}\alpha_{k}(a)\psi(f_{k})^{\frac{1}{2}})-\sum_{k\in K}\psi(f_{k})^{\frac{1}{2}}\alpha_{k}(a)\psi(f_{k})^{\frac{1}{2}}\|<\varepsilon_{0}.$
@@
    $\|a\psi(f_{k})-\psi(f_{k})a\|<\frac{\varepsilon_{0}}{|K|^{2}}\ \mbox{and}\ \|a\psi(f_{k})^{\frac{1}{2}}-\psi(f_{k})^{\frac{1}{2}}a\|<\frac{\varepsilon_{0}}{|K|^{2}}.$
@@
    If $k,k^{\prime}\in K$ satisfy $f_{k}f_{k^{\prime}}=0$, then
    $\|\psi(f_{k})^{\frac{1}{2}}\psi(f_{k^{\prime}})^{\frac{1}{2}}\|<\frac{\varepsilon_{0}}{|K|^{2}}.$
@@
    $(1-d-\varepsilon_{0})_{+}\precsim_{A^{\alpha}}y$.
@@
    $(1-d-\varepsilon_{0})_{+}\precsim_{A^{\alpha}}d$.
@@
    $\|dxd\|>1-\varepsilon_{0}$.
Denote the canonical conditional expectation by $E\colon A\to A^{\alpha}$. Define a completely positive contractive map $\varphi\colon A\to\overline{dA^{\alpha}d}$ by
for all $a\in F$. We will show $\varphi$ satisfies all properties. Clearly $\varphi(1)=d$, which is condition (1). Since $(1-d-\varepsilon)_{+}\leq(1-d-\varepsilon_{0})_{+}$, conditions (6), (7), (8), and (9) hold immediately.
For condition (2): Given $a,b\in F$, using 13 in the second step, (11), (12), (14), (15) in the third step, (10) and the fact that $(f_{k})_{k\in K}\subseteq C(G)$ is a partition of unity in the fourth step, and (13) in the fifth step, we have
@@
${}\approx_{2\varepsilon_{0}}\sum_{k^{\prime}\in K}\sum_{k\in K}\psi(f_{k^{\prime}})^{\frac{1}{2}}\alpha_{k^{\prime}}(a)\psi(f_{k^{\prime}})^{\frac{1}{2}}\psi(f_{k})^{\frac{1}{2}}\alpha_{k}(b)\psi(f_{k})^{\frac{1}{2}}$
${}\approx_{4\varepsilon_{0}}\sum_{k^{\prime}\in K}\sum_{k\in K}\psi(f_{k^{\prime}})\psi(f_{k})^{\frac{1}{2}}\alpha_{k}(ab)\psi(f_{k})^{\frac{1}{2}}$
@@
${}\approx_{\varepsilon_{0}}dE(\sum_{k\in K}\psi(f_{k})^{\frac{1}{2}}\alpha_{k}(ab)\psi(f_{k})^{\frac{1}{2}})$
@@
Therefore $\|\varphi(a)\varphi(b)-d\varphi(ab)\|<7\varepsilon_{0}=\varepsilon$, so condition (2) holds.
To prove condition (3): Let $a\in F$. Using (10) and the fact that $(f_{k})_{k\in K}\subseteq C(G)$ is a partition of unity in the first step, and (14) in the third step, we have
@@
${}<|K|\frac{\varepsilon_{0}}{|K|^{2}}<\varepsilon_{0}<\varepsilon.$
To verify condition (4): Let $a\in F_{2}\subseteq F$. Using (14) in the second step, (10) and the fact that $(f_{k})_{k\in K}\subseteq C(G)$ is a partition of unity in the fourth step, and (14) in the fifth step, we have
@@
${}\approx_{\varepsilon_{0}}E(\sum_{k\in K}\psi(f_{k})\alpha_{k}(a))$
@@
For condition (5): For every $n\in\mathbb{N}$, assume $\{S_{n}\}$ is an increasing sequence of finite subsets of $A$ such that $\cup_{n\in\mathbb{N}}S_{n}$ is dense in $A$ and $F\subseteq S_{n}$ for all $n\in\mathbb{N}$. For $\frac{1}{n}$, $S_{n}$, $F_{2}$, $x$, and $y$, there exist a contractive positive element $d_{n}\in A^{\alpha}$ and a completely positive contractive map $\varphi_{n}\colon A\to\overline{d_{n}A^{\alpha}d_{n}}$ such that
@@
    $\|\varphi_{n}(a)-d_{n}^{\frac{1}{2}}ad_{n}^{\frac{1}{2}}\|<\frac{1}{n}$ for all $a\in F_{2}$.
@@
Let $\pi_{A^{\alpha}}\colon l^{\infty}(\mathbb{N},A^{\alpha})\to(A^{\alpha})_{\infty}$ be the quotient map. Define a completely positive contractive order zero map $\Phi\colon A\to\overline{r(A^{\alpha})_{\infty}r}$ by $\Phi(a)=\pi_{A^{\alpha}}(\{\varphi_{1}(a),\varphi_{2}(a),\dots\})$ for all $a\in A$. Set $d=(d_{n})_{n=1}^{\infty}\in l^{\infty}(\mathbb{N},A^{\alpha})$ and $r=\pi_{A^{\alpha}}(d)$. We have
@@
From (34), $\|rxr\|=1$, we obtain $\|r\|=1$. Indeed, since $1=\|rxr\|\leq\|r\|\|x\|\|r\|=\|r\|^{2}\leq 1$, we have $\|r\|=1$.
Choose a finite set $T\subseteq F$ such that for every $a\in F$, there exists $b\in T$ with $\|b-a\|<\frac{\varepsilon}{3}$. By (34), we have $r\neq 0$. Therefore the order zero map $\Phi$ is nonzero. Since $A$ is simple, by [14, Proposition 5.3], $\|\Phi(a)\|=\|\Phi\|\|a\|$. Since $\|\Phi\|=\|\Phi(1)\|=\|r\|=1$, we deduce that the map is isometric. Therefore, for sufficiently large $n$ and all $b\in T$, we have
@@
Now, for sufficiently large $n$, set $\varphi=\varphi_{n}$ and $d=d_{n}$. Then conditions (1)-(9) hold. ∎
@@ 4. Simplicity of the fixed point algebra and the crossed product
For the case of finite groups, it follows from [12, Proposition 3.2] that the weak tracial Rokhlin property implies pointwise outerness. Therefore, by [32, Theorem 3.1], the crossed product is simple. However, without the discreteness of the group, pointwise outerness of the action does not imply simplicity of the crossed product.
@@
We first prove infinite-dimensionality. We will show that for any $n\in\mathbb{N}$, there exist nonzero mutually orthogonal positive elements $x_{1},\dots,x_{n}\in A^{\alpha}$. Since $A$ is simple and infinite-dimensional, there exist $y_{1},\dots,y_{n}\in A_{+}$ with $\|y_{k}\|=1$ for $k=1,\dots,n$. For given $\frac{1}{3}$, $n$, choose $0<\delta<\frac{1}{3}$ satisfying the condition of [33, Lemma 2.5.12].
@@
From (1), we have $\|\varphi(y_{j})\varphi(y_{k})\|<\delta$ when $j\neq k$. By [33, Lemma 2.5.12], there exist $x_{1},\dots,x_{n}\in(A^{\alpha})_{+}$ such that $\|x_{k}-\varphi(y_{k})\|<\frac{1}{3}$ for $k=1,\dots,n$ and $x_{j}x_{k}=0$ for $j\neq k$. Moreover, for $k=1,\dots,n$,
@@
[36, Lemma 3.5] In the setting of Definition 4.3, for any finite-dimensional unitary representation $\pi$ of $G$, the set $\overline{A_{2}(\pi)^{*}A_{2}(\pi)}$ is a closed two-sided ideal of $(B(\mathcal{H}_{\pi})\otimes A)^{\mathrm{Ad}(\pi)\otimes\alpha}$.
Recall ([38, Definition 7.1.4]) that an action of a compact group $G$ on a C*-algebra $A$ is saturated if a proper imprimitivity bimodule from $A^{\alpha}$ to $A\rtimes_{\alpha}G$ provides a Morita equivalence.
@@
[36, Theorem 3.6] Let $A$ be a C*-algebra, let $\alpha\colon G\to\mathrm{Aut}(A)$ be an action of a compact group on $A$. The following are equivalent:
@@
Set $d_{\pi}=\mathrm{dim}(\mathcal{H}_{\pi})$ and identify $\mathcal{H}_{\pi}$ with $\mathbb{C}^{d_{\pi}}$. Let $\delta=\min(\frac{1}{4d_{\pi}^{3}},\frac{1}{4})$. Equip $C(G)$ with the action $\mathrm{Lt}$. Denote by $\pi^{*}$ the function $g\mapsto\pi(g)^{*}$. A simple calculation shows that $\pi^{*}$ belongs to $C(G)_{2}(\pi)$. Indeed, fix $g\in G$, for $h\in G$, we have
@@
Therefore, $(\mathrm{id}_{B(\mathcal{H}_{\pi})}\otimes\mathrm{Lt}_{g})(\pi^{*})=\pi^{*}(\pi(g)\otimes 1_{C(G)})$, so $\pi^{*}\in C(G)_{2}(\pi)$. For $j,k=1,2,\dots,d_{\pi}$, let $\pi_{j,k}\in C(G)$ be the function $g\mapsto\pi(g)_{j,k}$. Applying Lemma 3.13, we obtain a contractive positive element $r\in(A_{\infty,\alpha}\cap A^{\prime})^{\alpha_{\infty}}$ and an equivariant completely positive contractive order zero map $\varphi\colon C(G)\to\overline{r(A_{\infty,\alpha}\cap A^{\prime})r}$ satisfying $\varphi(1)=r$ and $1-r\precsim_{p.w.-A^{\alpha}}r$. By Lemma 2.18 and [17, Corollary 2.10], we get an equivariant completely positive contractive order zero map $\sqrt{\varphi}$. Then

### v2

@@ 3. The weak tracial Rokhlin property with comparison
Without loss of generality, we can assume that $F$ is $\alpha$-invariant. For $F$ and $\frac{\varepsilon}{4}$, by Lemma 3.14, we can obtain $\delta>0$, a finite subset $K\subseteq G$, and a partition of unity $(f_{k})_{k\in K}\subseteq C(G)$ such that conditions (1) - (5) in Lemma 3.14 hold. Taking the continuous function $f(t)=t^{\frac{1}{2}}$, $t\in[0,1]$, and applying [5, Lemma 1.5] to $\frac{\varepsilon}{|K|^{2}}$, we obtain $\delta_{1}>0$. Let $\delta_{2}=\min(\frac{\varepsilon}{|K|^{2}},\frac{\varepsilon^{4}}{|K|^{8}},\delta_{1})$. Let $S=\{f_{k}\colon k\in K\}$, for the finite subset $S\subseteq C(G)$, the compact subset $F\subseteq A$, $\frac{1}{m}$, $x$ and $y$, by Lemma 3.6, we can obtain $d_{m}\in A^{\alpha}$ and a completely positive contractive equivariant $(S,F,\frac{1}{m})$-approximately central strongly order zero map $\psi_{m}\colon C(G)\to\overline{d_{m}Ad_{m}}$. Identify $C(G)\otimes A$ with $C(G,A)$, and for $m\in\mathbb{N}$, define a contractive linear map $\varphi_{m}\colon C(G,A)\to A$ by $\varphi_{m}(f\otimes a)=\psi_{m}(f)a$ for all $f\in C(G)$ and all $a\in A$. Clearly, $\varphi_{m}$ is equivariant, where we endow $C(G,A)$ with the diagonal action $\gamma$ of $G$. First, verify condition (5).
Given $a\in F$ and given $g\in G$, in the first step using the approximate centrality of $\psi_{m}$, in the second step using the equivariance of $\varphi_{m}$, and in the last step applying condition (5) from the conclusion of Lemma 3.14 and noting that for the averaging process $\gamma_{g}(\varphi(a))=\varphi(a)$, we have
@@
$\leq$ ${}\limsup_{m\to\infty}\|\gamma_{g}(\sum_{k\in K}f_{k}\otimes\alpha_{k}(a))-\sum_{k\in K}f_{k}\otimes\alpha_{k}(a)\|<\frac{\varepsilon}{2}.$
Choose a sufficiently large $m>\frac{1}{\delta_{2}}$ such that
@@
Then for such $m$, condition (1) is obvious. Due to the continuity of $\alpha$, condition (2) is also obvious. Conditions (3) and (4) come from Lemma 3.14. Conditions (8) - (11) come from Lemma 3.6.
For condition (6), for all $a\in F$ and all $k\in K$, we have
$\|a\psi(f_{k})-\psi(f_{k})a\|<\delta_{2}=\min(\frac{\varepsilon}{|K|^{2}},\frac{\varepsilon^{4}}{|K|^{8}},\ \delta_{1}).$
Therefore, by [5, Lemma 1.5], for all $a\in F$ and all $k\in K$, we have
$\|a\psi(f_{k})^{\frac{1}{2}}-\psi(f_{k})^{\frac{1}{2}}a\|<\frac{\varepsilon}{|K|^{2}}.$
For condition (7), whenever $k,k^{\prime}\in K$ satisfy $f_{k}f_{k^{\prime}}=0$, we have
$\|\psi(f_{k})\psi(f_{k^{\prime}})\|<\delta_{2}\leq\frac{\varepsilon}{|K|^{2}}.$
Hence,
$\|\psi(f_{k})^{\frac{1}{2}}\psi(f_{k^{\prime}})^{\frac{1}{2}}\|^{4}=$ ${}\|\psi(f_{k})^{\frac{1}{2}}\psi(f_{k^{\prime}})\psi(f_{k})\psi(f_{k^{\prime}})\psi(f_{k})^{\frac{1}{2}}\|$
$\leq$ ${}\|\psi(f_{k})\|\|\psi(f_{k^{\prime}})\|\|\psi(f_{k})\psi(f_{k^{\prime}})\|$
$\leq$ ${}\|\psi(f_{k})\psi(f_{k^{\prime}})\|<\delta_{2}.$
Thus $\|\psi(f_{k})^{\frac{1}{2}}\psi(f_{k^{\prime}})^{\frac{1}{2}}\|<\delta_{2}^{\frac{1}{4}}\leq\frac{\varepsilon}{|K|^{2}}$. ∎
We now present the most important theorem of this section, which is a key tool for transferring properties from the original algebra to the fixed point algebra. The proof references [19, Theorem 3.1].
@@
Let $\varepsilon_{0}=\frac{\varepsilon}{7}$, take the continuous function $f(t)=t^{\frac{1}{2}}$, $t\in[0,1]$, and apply [5, Lemma 1.5] to $\varepsilon_{0}$ to obtain $\delta_{1}>0$. Let $\delta_{2}=\min(\varepsilon_{0},\ \delta_{1})$. Without loss of generality, we may assume that $\|a\|\leq 1$ for all $a\in F_{1}\cup F_{2}$. For the compact set $F=F_{1}\cup F_{2}$ (without loss of generality, assume it is $\alpha$-invariant), $\delta_{2}$, $x\in A_{+}$ with $\|x\|=1$, and $y\in(A^{\alpha})_{+}\setminus\{0\}$, using Proposition 3.15, we can obtain a positive number $\delta>0$, a finite subset $K\subseteq G$, a partition of unity $(f_{k})_{k\in K}\subseteq C(G)$, a contractive positive element $d\in A^{\alpha}$, and a completely positive contractive map $\psi\colon C(G)\to\overline{dAd}$ such that
@@
    If $g,g^{\prime}\in G$ and $d(g,g^{\prime})<\delta$, then for all $a\in F$, $\|\alpha_{g}(a)-\alpha_{g^{\prime}}(a)\|<\frac{\delta_{2}}{|K|^{2}}$.
@@
    Whenever $k,k^{\prime}\in K$ satisfy $f_{k}f_{k^{\prime}}\neq 0$, then $d(k,k^{\prime})<\delta$.
@@
    $\|\alpha_{g}(\sum_{k\in K}\psi(f_{k})^{\frac{1}{2}}\alpha_{k}(a)\psi(f_{k})^{\frac{1}{2}})-\sum_{k\in K}\psi(f_{k})^{\frac{1}{2}}\alpha_{k}(a)\psi(f_{k})^{\frac{1}{2}}\|<\delta_{2}.$
@@
    $\|a\psi(f_{k})-\psi(f_{k})a\|<\frac{\delta_{2}}{|K|^{2}}\ \mbox{and}\ \|a\psi(f_{k})^{\frac{1}{2}}-\psi(f_{k})^{\frac{1}{2}}a\|<\frac{\delta_{2}}{|K|^{2}}.$
@@
    Whenever $k,k^{\prime}\in K$ satisfy $f_{k}f_{k^{\prime}}=0$, we have
    $\|\psi(f_{k})^{\frac{1}{2}}\psi(f_{k^{\prime}})^{\frac{1}{2}}\|<\frac{\delta_{2}}{|K|^{2}}.$
@@
    $(1-d-\delta_{2})_{+}\precsim_{A^{\alpha}}y$.
@@
    $(1-d-\delta_{2})_{+}\precsim_{A^{\alpha}}d$.
@@
    $\|dxd\|>1-\delta_{2}$.
Denote the canonical conditional expectation by $E\colon A\to A^{\alpha}$. Define a completely positive contractive map $\varphi\colon A\to\overline{dA^{\alpha}d}$ by, for all $a\in F$,
$\varphi(a)=E(\sum_{k\in K}\psi(f_{k})^{\frac{1}{2}}\alpha_{k}(a)\psi(f_{k})^{\frac{1}{2}}).$
We will show that $\varphi$ satisfies all the properties. Clearly $\varphi(1)=d$, which is condition (1). Since $(1-d-\varepsilon)_{+}\leq(1-d-\delta_{2})_{+}$, conditions (6), (7), (8), and (9) immediately hold.
For condition (2), given $a,b\in F$, using (13) in the second step, (11), (12), (14), (15) in the third step, (10) and the fact that $(f_{k})_{k\in K}\subseteq C(G)$ is a partition of unity in the fourth step, and (13) in the fifth step, we have
@@
${}\approx_{2\delta_{2}}\sum_{k^{\prime}\in K}\sum_{k\in K}\psi(f_{k^{\prime}})^{\frac{1}{2}}\alpha_{k^{\prime}}(a)\psi(f_{k^{\prime}})^{\frac{1}{2}}\psi(f_{k})^{\frac{1}{2}}\alpha_{k}(b)\psi(f_{k})^{\frac{1}{2}}$
${}\approx_{4\delta_{2}}\sum_{k^{\prime}\in K}\sum_{k\in K}\psi(f_{k^{\prime}})\psi(f_{k})^{\frac{1}{2}}\alpha_{k}(ab)\psi(f_{k})^{\frac{1}{2}}$
@@
${}\approx_{\delta_{2}}dE(\sum_{k\in K}\psi(f_{k})^{\frac{1}{2}}\alpha_{k}(ab)\psi(f_{k})^{\frac{1}{2}})$
@@
Hence $\|\varphi(a)\varphi(b)-d\varphi(ab)\|<7\delta_{2}\leq\varepsilon$, i.e., condition (2) holds.
To prove condition (3), let $a\in F$. Using (10) and the fact that $(f_{k})_{k\in K}\subseteq C(G)$ is a partition of unity in the first step, and (14) in the third step, we have
@@
${}<|K|\frac{\delta_{2}}{|K|^{2}}<\delta_{2}=\min(\varepsilon_{0},\ \delta_{1})<\varepsilon.$
Then by [5, Lemma 1.5], we have
(3.1) $\|d^{\frac{1}{2}}a-ad^{\frac{1}{2}}\|<\varepsilon_{0}.$
To verify condition (4), let $a\in F_{2}\subseteq F$. Using (14) in the second step, (10) and the fact that $(f_{k})_{k\in K}\subseteq C(G)$ is a partition of unity in the fourth step, and (3.1) in the fifth step, we have
@@
${}\approx_{\delta_{2}}E(\sum_{k\in K}\psi(f_{k})\alpha_{k}(a))$
@@
For condition (5), for any $n\in\mathbb{N}$, assume $\{S_{n}\}$ is a sequence of finite subsets in $A$ such that $\cup_{n\in\mathbb{N}}S_{n}$ is dense in $A$ and $F\subseteq S_{n}$ for all $n\in\mathbb{N}$. For $\frac{1}{n}$, $S_{n}$, $F_{2}$, $x$, and $y$, there exist a contractive positive element $d_{n}\in A^{\alpha}$ and a completely positive contractive map $\varphi_{n}\colon A\to\overline{d_{n}A^{\alpha}d_{n}}$ such that
@@
    For all $a\in F_{2}$, $\|\varphi_{n}(a)-d_{n}^{\frac{1}{2}}ad_{n}^{\frac{1}{2}}\|<\frac{1}{n}$.
@@
Let $\pi_{A^{\alpha}}\colon l^{\infty}(\mathbb{N},A^{\alpha})\to(A^{\alpha})_{\infty}$ be the quotient map. Define a completely positive contractive order zero map $\Phi\colon A\to\overline{r(A^{\alpha})_{\infty}r}$ by $\Phi(a)=\pi_{A^{\alpha}}(\{\varphi_{1}(a),\varphi_{2}(a),\dots\})$ for all $a\in A$. Denote $d=(d_{n})_{n=1}^{\infty}\in l^{\infty}(\mathbb{N},A^{\alpha})$ and $r=\pi_{A^{\alpha}}(d)$. We have
@@
From $\|rxr\|=1$ in (34), we can deduce $\|r\|=1$. Indeed, since $\|rxr\|=1$, we have $1=\|rxr\|\leq\|r\|\|x\|\|r\|=\|r\|^{2}\leq 1$, so $\|r\|=1$.
Choose a finite set $T\subseteq F$ such that for every $a\in F$ there exists $b\in T$ with $\|b-a\|<\frac{\varepsilon}{3}$. From (34), we have $r\neq 0$. Therefore, the order zero map $\Phi$ is nonzero. Since $A$ is simple, by [15, Proposition 5.3], $\|\Phi(a)\|=\|\Phi\|\|a\|$. Because $\|\Phi\|=\|\Phi(1)\|=\|r\|=1$, we conclude that this map is isometric. Hence, for sufficiently large $n$ and all $b\in T$, we have
@@
Now, for sufficiently large $n$, let $\varphi=\varphi_{n}$ and $d=d_{n}$, then conditions (1) - (9) hold. ∎
@@ 4. Simplicity of the fixed point algebra and the crossed product
For the case of finite groups, it follows from [13, Proposition 3.2] that the weak tracial Rokhlin property implies pointwise outerness. Therefore, by [33, Theorem 3.1], the crossed product is simple. However, without the discreteness of the group, pointwise outerness of the action does not imply simplicity of the crossed product.
@@
We first prove infinite-dimensionality. We will show that for any $n\in\mathbb{N}$, there exist nonzero mutually orthogonal positive elements $x_{1},\dots,x_{n}\in A^{\alpha}$. Since $A$ is simple and infinite-dimensional, there exist $y_{1},\dots,y_{n}\in A_{+}$ with $\|y_{k}\|=1$ for $k=1,\dots,n$. For given $\frac{1}{3}$, $n$, choose $0<\delta<\frac{1}{3}$ satisfying the condition of [34, Lemma 2.5.12].
@@
From (1), we have $\|\varphi(y_{j})\varphi(y_{k})\|<\delta$ when $j\neq k$. By [34, Lemma 2.5.12], there exist $x_{1},\dots,x_{n}\in(A^{\alpha})_{+}$ such that $\|x_{k}-\varphi(y_{k})\|<\frac{1}{3}$ for $k=1,\dots,n$ and $x_{j}x_{k}=0$ for $j\neq k$. Moreover, for $k=1,\dots,n$,
@@
[37, Lemma 3.5] In the setting of Definition 4.3, for any finite-dimensional unitary representation $\pi$ of $G$, the set $\overline{A_{2}(\pi)^{*}A_{2}(\pi)}$ is a closed two-sided ideal of $(B(\mathcal{H}_{\pi})\otimes A)^{\mathrm{Ad}(\pi)\otimes\alpha}$.
Recall ([40, Definition 7.1.4]) that an action of a compact group $G$ on a C*-algebra $A$ is saturated if a proper imprimitivity bimodule from $A^{\alpha}$ to $A\rtimes_{\alpha}G$ provides a Morita equivalence.
@@
[37, Theorem 3.6] Let $A$ be a C*-algebra, let $\alpha\colon G\to\mathrm{Aut}(A)$ be an action of a compact group on $A$. The following are equivalent:
@@
Below we give an equivariant version of Lemma 2.18, that is, an equivariant completely positive order zero map remains an equivariant completely positive order zero map after continuous functional calculus by a nonnegative function.
Lemma 4.6.
Let $G$ be a locally compact group, and let $A$ be a unital C*-algebra, $B$ a C*-algebra. Let $\alpha\colon G\to\mathrm{Aut}(A)$ and $\gamma\colon G\to\mathrm{Aut}(B)$ be actions of $G$ on $A$ and $B$, respectively. Let $\varphi\colon A\to B$ be a completely positive contractive order zero map, and let $f\in C_{0}((0,1])$ be a positive function. Let $C,h$, and $\pi_{\varphi}$ be as in Theorem 2.16. Then, the map defined by
$f(\varphi)(a)\colon=f(h)\pi_{\varphi}(a)\quad\forall a\in A,$
$f(\varphi)\colon A\to C\subset B$
is a well-defined equivariant completely positive order zero map. If the norm of $f$ is at most one, then $f(\varphi)$ is also contractive.
Proof.
Here we only verify the equivariance. Since $\varphi$ is equivariant, for any $g\in G$ we have $\varphi(\alpha_{g}(1))=\beta_{g}(\varphi(1))$, i.e., $\varphi(1)=\beta_{g}(\varphi(1))$. By the arbitrariness of $g$, it follows that $\varphi(1)$ is $\beta$-invariant. Then for any $x\in A$, $g\in G$, by Lemma 2.18 we have
$\pi_{\varphi}(\alpha_{g}(x))h=$ ${}\varphi(\alpha_{g}(x))=\gamma_{g}(\varphi(x))=\gamma_{g}(\pi_{\varphi}(x)h)=\gamma_{g}(\pi_{\varphi}(x))h.$
Hence $(\pi_{\varphi}(\alpha_{g}(x))-\gamma_{g}(\pi_{\varphi}(x)))h=0$. From the continuous functional calculus for positive elements, it is easy to see that $(\pi_{\varphi}(\alpha_{g}(x))-\gamma_{g}(\pi_{\varphi}(x)))f(h)=0$, and $f(h)$ is also $\beta$-invariant.
Therefore, we have $\pi_{\varphi}(\alpha_{g}(x))f(h)=\gamma_{g}(\pi_{\varphi}(x))f(h)$, i.e., $f(\varphi)(\alpha_{g}(x))=\gamma_{g}(f(\varphi)(x))$. ∎
@@
Let $d_{\pi}=\mathrm{dim}(\mathcal{H}_{\pi})$ and identify $\mathcal{H}_{\pi}$ with $\mathbb{C}^{d_{\pi}}$. Let $\delta=\min(\frac{1}{4d_{\pi}^{3}},\frac{1}{4})$. Endow $C(G)$ with the action $\mathrm{Lt}$. Denote by $\pi^{*}$ the function $g\mapsto\pi(g)^{*}$. A simple calculation shows that $\pi^{*}$ belongs to $C(G)_{2}(\pi)$. Indeed, fix $g\in G$, for $h\in G$, we have
@@
Thus, $(\mathrm{id}_{B(\mathcal{H}_{\pi})}\otimes\mathrm{Lt}_{g})(\pi^{*})=\pi^{*}(\pi(g)\otimes 1_{C(G)})$, so $\pi^{*}\in C(G)_{2}(\pi)$. For $j,k=1,2,\dots,d_{\pi}$, let $\pi_{j,k}\in C(G)$ be the function $g\mapsto\pi(g)_{j,k}$. Applying Lemma 3.13, we obtain a contractive positive element $r\in(A_{\infty,\alpha}\cap A^{\prime})^{\alpha_{\infty}}$ and an equivariant completely positive contractive order zero map $\varphi\colon C(G)\to\overline{r(A_{\infty,\alpha}\cap A^{\prime})r}$ satisfying $\varphi(1)=r$ and $1-r\precsim_{p.w.-A^{\alpha}}r$. Note that $r$ is $\alpha_{\infty}$-invariant. By Lemma 4.6, it is easy to see that we obtain an equivariant completely positive contractive order zero map $\sqrt{\varphi}$. Then